OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates. (-5, 5pi/3)

OpenStudy (anonymous):

-(5/2), 5sqrt3/2

OpenStudy (anonymous):

\[(-5\cos(\frac{5\pi}{3}),-5\sin(\frac{5\pi}{3}))\]